# Exploring Imbalanced Annotations for Effective In-Context Learning

### v1

@@ A.6.5 Can our method improve real-world imbalanced dataset?
Imbalance Ratio 0 10 50 100 0 10 50 100
Random 12.07$\pm$0.25 15.73$\pm$0.34 20.27$\pm$1.25 21.33$\pm$0.52 12.87$\pm$0.47 20.33$\pm$0.96 22.87$\pm$2.07 23.87$\pm$0.41
+Ours 11.27$\pm$0.66 20.13$\pm$0.98 19.13$\pm$1.11 22.40$\pm$2.41 13.20$\pm$0.28 15.60$\pm$0.86 20.40$\pm$1.14 22.80$\pm$0.43
TopK 31.40$\pm$0.00 29.67$\pm$1.20 25.87$\pm$0.68 25.33$\pm$0.66 36.20$\pm$0.16 32.47$\pm$1.06 27.40$\pm$0.99 25.47$\pm$0.09
+Ours 30.80$\pm$0.00 29.83$\pm$1.01 27.87$\pm$1.05 26.20$\pm$1.23 33.60$\pm$0.85 31.17$\pm$0.29 29.93$\pm$0.52 27.13$\pm$0.98
DPP 31.93$\pm$0.09 30.00$\pm$0.49 27.47$\pm$1.04 26.20$\pm$1.02 38.33$\pm$0.25 32.93$\pm$1.11 28.07$\pm$0.93 26.53$\pm$0.25
+Ours 31.60$\pm$0.85 30.40$\pm$1.57 27.93$\pm$0.93 27.07$\pm$2.10 36.00$\pm$0.00 30.73$\pm$0.66 29.80$\pm$0.71 30.03$\pm$0.73
VoteK 28.80$\pm$0.16 29.60$\pm$0.59 27.07$\pm$1.39 25.73$\pm$0.41 36.13$\pm$0.09 32.47$\pm$1.06 27.40$\pm$0.99 25.47$\pm$0.09
+Ours 28.93$\pm$0.74 29.13$\pm$1.48 27.87$\pm$0.62 26.47$\pm$0.66 36.07$\pm$0.34 32.47$\pm$1.06 29.27$\pm$0.57 28.33$\pm$0.82
ConE 28.47$\pm$1.32 27.60$\pm$1.66 24.93$\pm$1.05 23.87$\pm$0.93 36.87$\pm$0.38 29.93$\pm$0.52 25.93$\pm$0.09 24.13$\pm$0.47
+Ours 29.13$\pm$0.38 27.93$\pm$0.41 26.33$\pm$1.72 25.20$\pm$0.49 34.53$\pm$0.09 31.07$\pm$0.19 27.60$\pm$1.23 27.47$\pm$0.47
ByDC 31.73$\pm$0.09 29.27$\pm$0.84 26.80$\pm$1.02 26.33$\pm$0.84 37.53$\pm$0.34 30.87$\pm$0.62 26.47$\pm$0.57 24.80$\pm$0.43
+Ours 31.47$\pm$0.25 30.27$\pm$0.19 27.80$\pm$0.72 27.27$\pm$0.74 35.20$\pm$0.16 31.53$\pm$0.34 28.80$\pm$1.50 28.13$\pm$1.46
Baselines 27.40$\pm$0.32 26.98$\pm$0.85 25.40$\pm$1.07 24.80$\pm$0.73 32.99$\pm$0.28 29.83$\pm$0.89 26.36$\pm$0.94 25.05$\pm$0.29
+Ours 27.20$\pm$0.48 27.95$\pm$0.94 26.16$\pm$1.02 25.77$\pm$1.27 31.43$\pm$0.29 28.76$\pm$0.57 27.63$\pm$0.95 27.32$\pm$0.82
Table 8: The average Rouge score of CodeSearchNet dataset across six selection methods with various imbalance ratios. Bold numbers are the best results.
Imbalance Ratio 0 10 50 100
Baselines 15.99 13.26 11.41 9.84
+Over-Sampling 16.12 13.82 10.07 8.56
+Under-Sampling 16.08 10.17 10.06 9.20
+Stratified-Sampling 10.59 9.95 9.54 9.24
+Re-Weighting 15.86 13.33 8.59 7.96
+Ours 15.71 13.64 11.86 10.90
Figure 18: Average accuracy of Emotion tasks across six existing selection methods using OPT-6.7B (Zhang et al., 2022a).

### v2

@@ E.7 Main results for each dataset
Imbalance Ratio 1 10 50 100 1 10 50 100
Random 42.63$\pm$0.38 39.80$\pm$0.93 37.60$\pm$0.73 37.20$\pm$0.13 44.27$\pm$0.36 42.10$\pm$0.67 38.90$\pm$0.27 38.17$\pm$0.22
+Ours 42.90$\pm$0.80 42.97$\pm$1.42 41.40$\pm$1.07 42.20$\pm$0.13 45.40$\pm$0.60 45.63$\pm$0.36 44.40$\pm$1.73 43.30$\pm$1.40
TopK 47.30$\pm$0.00 45.73$\pm$0.58 41.13$\pm$0.64 40.27$\pm$1.36 49.70$\pm$0.00 46.97$\pm$0.38 43.30$\pm$0.27 42.17$\pm$0.49
+Ours 46.97$\pm$0.44 45.40$\pm$0.67 43.93$\pm$0.89 42.07$\pm$0.64 48.10$\pm$0.00 46.60$\pm$0.40 44.93$\pm$0.11 43.83$\pm$0.36
DPP 46.80$\pm$0.00 44.83$\pm$0.82 41.97$\pm$0.84 40.57$\pm$0.44 48.10$\pm$0.00 46.33$\pm$0.89 43.90$\pm$0.13 42.77$\pm$0.18
+Ours 46.27$\pm$0.71 44.60$\pm$0.33 42.37$\pm$0.91 41.97$\pm$0.36 48.13$\pm$0.04 45.83$\pm$0.78 44.20$\pm$0.47 44.03$\pm$0.49
VoteK 48.10$\pm$0.20 45.60$\pm$0.33 41.77$\pm$0.36 40.73$\pm$0.49 49.20$\pm$0.00 46.70$\pm$0.47 43.40$\pm$0.53 42.07$\pm$0.56
+Ours 47.77$\pm$0.31 45.40$\pm$0.53 42.60$\pm$0.80 41.63$\pm$0.31 49.00$\pm$0.13 46.73$\pm$0.16 44.10$\pm$0.53 43.17$\pm$0.58
ConE 46.20$\pm$0.00 44.17$\pm$0.69 41.23$\pm$1.18 39.70$\pm$0.47 48.63$\pm$0.04 46.40$\pm$0.53 42.73$\pm$0.56 41.87$\pm$0.29
+Ours 46.53$\pm$0.44 44.50$\pm$0.73 42.57$\pm$0.58 40.87$\pm$0.64 48.43$\pm$0.11 46.57$\pm$0.42 43.23$\pm$0.22 42.53$\pm$0.38
ByDC 47.70$\pm$0.33 45.27$\pm$0.16 41.93$\pm$0.58 40.80$\pm$0.53 48.97$\pm$0.09 46.37$\pm$0.36 43.07$\pm$0.11 42.00$\pm$0.33
+Ours 47.43$\pm$0.18 45.60$\pm$0.40 42.27$\pm$0.51 42.10$\pm$0.20 48.37$\pm$0.09 46.33$\pm$0.24 43.57$\pm$0.29 43.57$\pm$0.29
Vanilla 46.46$\pm$0.15 44.23$\pm$0.59 40.94$\pm$0.72 39.88$\pm$0.57 48.14$\pm$0.08 45.81$\pm$0.55 42.55$\pm$0.31 41.51$\pm$0.34
+Ours 46.31$\pm$0.48 44.74$\pm$0.68 42.52$\pm$0.79 41.81$\pm$0.38 47.91$\pm$0.16 46.28$\pm$0.39 44.07$\pm$0.56 43.46$\pm$0.58
@@
Figure 6: The Kullback-Leibler (KL) divergence between $P_{c}(X|Y)$ and $P_{t}(X|Y)$ for each class in the Yahoo datasets. The KL divergence validates the existence of conditional bias in imbalanced datasets.
(a) Amazon
(b) Yelp
(c) Yahoo
Figure 7: The comparison of average accuracy for each class in the Amazon (a), Yelp (b) and Yahoo (c) datasets across various imbalance ratios.